OpenStudy (anonymous):

find the height of a flag pole if the angle of elevation of its top changes from 30 to 45 degrees as the observer advances 10 ft. towards the base

OpenStudy (pawanyadav):

Let height be h Then distance from pole=h/tan30 =h√3. (Tan30=1/√3) Again Angle is 45 Than distance =h/tan45 =h. (Tan45=1) Now According to question h√3-h=10. Solve for h now.

OpenStudy (pawanyadav):

Make diagram for more illustration Remember tanx=perpendicular/base So. Length of base=perpendicular/tanx

OpenStudy (anonymous):

why did you use \[\sqrt{3}\] ?

OpenStudy (anonymous):

@Pawanyadav

OpenStudy (pawanyadav):

tan30. is what ?

OpenStudy (anonymous):

\[\sqrt{3}/3\]

OpenStudy (pawanyadav):

We can write it 1/√3

OpenStudy (anonymous):

oh i see thank you :)

OpenStudy (pawanyadav):

As 3=√3×√3 Canceling one √3 we get √3/3=1/√3